OpenStudy (anonymous):

Find the result when 240 is increased by 5%

OpenStudy (anonymous):

i think it is 240x 0.5 which equals 120

OpenStudy (anonymous):

oh wait scratch that!! that is not right

OpenStudy (anonymous):

240+12=252

OpenStudy (anonymous):

x/240 * 100=5 x=5*240/100 x=12 12 is 5% of 240 so 240+12=252